Leading men. Bob hope was born 2 years after Cark Gable and 2 years before Henry Fonda. In 1951, after all three of the had their birthdays, the sum of their ages was 144. In what years were they born?



Answer by Edwin McCravy(20086)

Leading men. Bob hope was born 2 years after Cark Gable and 2 years before Henry Fonda. In 1951, after all three of the had their birthdays, the sum of their ages was 144. In what years were they born?

Let Bob Hope's birth year be N

>>...Bob hope was born 2 years after Clark Gable...<<

Therefore Clark Gable's birth year is N-2

>>...Bob hope was born...2 years before Henry Fonda...<<

Therefore Henry Fonda's birth year is N+2.

In 1951, Bob Hope was 1951 - N

in 1951, Clark Gable was 1951 - (N-2) 

1n 1951, Henry Fonda was 1951 - (N+2)

>>...In 1951, after all three of the had their birthdays, 
the sum of their ages was 144...<<

Therefore 


1951 - N + 1951 - (N-2) + 1951 - (N+2) = 144

1951 - N + 1951 - N + 2 + 1951 - N - 2 = 144

                             5853 - 3N = 144

                                   -3N = -5709
  
                                     N = 1903

So Bob Hope was born in 1903.    

>>...Bob hope was born 2 years after Clark Gable...<<

So Clark Gable was born in 1901.

>>...Bob hope was born...2 years before Henry Fonda...<<

So Henry Fonda was born in 1905.

Edwin